# Retrieval of fundamental material parameters of monolayer transition metal dichalcogenides from experimental exciton energies: An analytical approach

### v1

@@ III Exciton reduced mass from magnetoexciton energies
where $\alpha=r_{0}/(\kappa a_{0})$; $r=\sqrt{x^{2}+y^{2}}$; energy $\varepsilon$ and coordinates $x,y$ are given in the Rydberg unit Ry and Bohr radius $a_{0}$, respectively; B is dimensionless magnetic intensity measured in the unit of $B_{0}=(m_{e}/\hbar e)\text{Ry}$; $\hbar$ and $e$ are the reduced Planck constant and electron electric charge, respectively. Here, in Eq. (III), we use the Laplace transform for the Rytova-Keldysh potential describing the electron-hole interaction, given in Ref.  [32].
The Schrödinger equation (III) contains three material parameters: $\kappa$, $r_{0}$, and $\mu$. However, we can exclude two of them, $\kappa$ and $r_{0}$, from the equation using formulas (10) and (11). Consequently, the equation consists of only one material parameter, $\mu$. Previous studies show the strong dependence of the magnetoexciton energies on the exciton reduced mass. Therefore, we consider $\mu$ a variable parameter and vary it so that the calculated magnetoexciton energies $\varepsilon_{ns}(B),\,(n=1,2,3,4)$ best fit with the experimental data $E_{ns}(B)-E_{g}$, where the bandgap energy $E_{g}$, in turn, can be calculated by formula (9) from the experimental data. This way, we retrieve the exciton reduced mass $\mu$ from the experimental magnetoexciton energies $E_{ns}(B)$. We note that a numerical method based on the Feranchuk-Komarov method was proposed in Ref. [32] to solve the Schrödinger like Eq. (III) with the codes available upon request. This method is speedy and gives highly accurate magnetoexciton energies.
For demonstration purposes, in Fig. 3 (a), we plot the magnetoexciton energy spectrum for monolayer WSe2 on the hBN substrate calculated with different values of $\mu$ (0.17, 0.18, and 0.19 $m_{e}$). Other material parameters ($E_{g}$, $r_{0}$, $\kappa$) are obtained by Eqs. (9), (10), and (11) with the experimental exciton energies ($E_{1s}$, $E_{2s}$, $E_{3s}$) of Ref. [29]. We also plot the experimental magnetoexciton energies from Ref. [29] for comparison. As discussed above, by adjusting $\mu$ so that the theoretical model agrees with experimental data, we obtain the best-fit values of the exciton reduced mass, $\mu=0.18m_{e}$. Other figures, Figs. 3 (b) and (c), are plotted in the same manner but for experimental data of Refs. [27] and [37].
Figure 3: Magnetoexciton energies in monolayer WSe2 encapsulated by hBN slabs, calculated by the numerical method (lines) compared with experimental data (blue triangles) of Refs. [29] (a), [27] (b), and [37] (c). Bandgap energy $E_{g}$ and screening length $r_{0}$ are retrieved from the experimental data of these references. Exciton reduced masses for these cases are the best-fit values (black line), which leads to the surrounding dielectric constants: $\kappa=4.41$ (a), $4.26$ (b), and $3.35$ (c).
@@
From the above-formulated analytical expressions, we arrive at a practical scheme, shown in Fig. 1, to retrieve material parameters from experimental magnetoexciton energies.
@@
Applying this retrieval scheme to the experimental data of Refs. [27, 29, 37, 36, 38], we retrieve the material parameters for monolayer WSe2. We provide them in Table 2 and compare them with the ones presented in these references. Knowing the material parameters, we also calculate the other essential properties, such as the diamagnetic constants ($\sigma_{1s}$, $\sigma_{2s}$) and exciton radii ($r_{1s}$, $r_{2s}$). More data retrieved in the same way from Ref. [28] for other TMDC monolayers, such as WS2, MoS2, and MoSe2, are also given in Table 2.
@@
The first remark from Table 2 is that although various experiments were performed for the same material, monolayer WSe2 encapsulated by hBN slabs, the surrounding dielectric constants $\kappa$ from these experiments differ due to variations in the experimental setups. However, the exciton reduced mass $\mu$ is weakly dependent on the surrounding dielectric constant $\kappa$, as shown in Table 2. For various experiments, it almost equals $0.19\,m_{e}$. With the exciton reduced mass $\mu$ known, our method is also applicable to photoluminescence experiments without a magnetic field, as shown in Table 2 for Ref. [38]. Indeed, we can determine the bandgap energy $E_{g}$ and screening length $r_{0}$ using Eqs. (9) and (10) within the experimental exciton energies $E_{1s}$, $E_{2s}$, and $E_{3s}$. The surrounding dielectric constant $\kappa$ is then calculated by Eq. (11), putting the exciton reduced mass’s value $\mu=0.19\,m_{e}$, taken from other experiments. Moreover, the surrounding dielectric constants for different experiments are nearly identical (except in Chen et al. [37]), with values approximately equal to the bulk h-BN dielectric constant of 4.5, as referenced in [5].
Table 1: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WSe2 retrieved from magnetoexciton energies of different experiments [27, 29, 37, 36, 38] using analytical expressions (9), (10), and (11) and compared with those given in these references.
WSe2 $r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$ $\sigma_{1s}$ $r_{1s}$ $\sigma_{2s}$ $r_{2s}$
(nm) $(m_{e})$ (eV) (meV) (meV) (meV) ($\mu$eV/T2) (nm) ($\mu$eV/T2) (nm)
Stier et al. [27] – exp. (169) 130 22 0.31$\pm$0.02 1.7$\pm$0.1 4.6$\pm$0.2 6.6$\pm$0.4
[27] – theory 4.50 0.20 4.50 1.890 167 124 21.3 1.67 6.96
Present work’s theory 4.34 0.19 4.26 1.892 168.6 129.9 22.0 0.33 1.69 5.69 7.01
Liu et al. [29] – exp. (167) 131 21 0.24$\pm$0.1 1.6$\pm$0.4 6.4 $\pm$0.2 $8.24\pm$0.13
[29] – theory 5.00 0.20 3.97 1.884 172 0.31 4.86
Present work’s theory 3.89 0.18 4.41 1.879 167.2 130.9 21.0 0.36 1.71 6.46 7.27
Chen et al. [37] – exp. (177) 131 26 0.5 2.2 5.8 7.6
[37] – theory 4.51 0.22 4.50 1.900 170 128 23 0.25 1.6 4.18 6.5
Present work’s theory 5.52 0.19 3.35 1.907 179.9 130.9 26.0 0.34 1.72 4.89 6.5
Molas et al. [36] – exp. (167) 129 21.5
[36] – theory 4.50 0.20 4.50 1.873 167 128.0 21.2
Present work 4.25 0.19 4.35 1.873 166.8 129.1 21.5 0.33 1.69 5.73 7.04
Takahashi et al. [38] – exp. (167) 129.1 21.6
[38] – theory 3.45 0.20 5.01 1.862 164
Present work 4.28 0.19* 4.33 1.865 167.0 129.1 21.6 0.33 1.69 5.72 7.03
Table 2: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WS2, MoSe2, and MoS2 retrieved from magnetoexciton energies of the experiment [28] using analytical expressions (9), (10), and (11) and compared with those given in this reference.
$r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$ $\sigma_{1s}$ $r_{1s}$ $\sigma_{2s}$ $r_{2s}$
(nm) $(m_{e})$ (eV) (meV) (meV) (meV) ($\mu$eV/T2) (nm) ($\mu$eV/T2) (nm)
WS2
Goryca et al. [28] – exp. (181) 141 23
[28] – theory 3.40 0.175 4.35 2.238 180 1.8
Present work’s theory 3.74 0.175 4.10 2.239 180.9 140.9 23.0 0.36 1.69 6.33 7.10
MoSe2
Goryca et al. [28] – exp. (232) 171 33
[28] – theory 3.90 0.35 4.40 1.874 231 0.07 1.1
Present work’s theory 4.03 0.355 4.28 1.875 231.6 171.1 32.8 0.08 1.1 1.07 4.15
MoS2
Goryca et al. [28] – exp. (221) 170 29
[28] – theory 3.40 0.27 4.40 2.161 222 0.12 1.2
Present work’s theory 3.34 0.275 4.46 2.160 220.9 170.0 28.9 0.12 1.23 2.06 5.07
@@
The second remark from Tables 2 and 2 is the slight discrepancy between the bandgap energies calculated by our method and those in the references. By applying the bandgap energies we calculated, the experimental binding energies for the $1s$ state (in the parentheses) coincide with the prediction of our theory. The consistency of our predicted energy transitions $\Delta E_{1s-2s}$ and $\Delta E_{2s-3s}$ to the experimental data is much better than the theoretical modeling by the Keldysh potential in Refs. [27, 29, 37] or by the Kratzer potential in Ref. [36] for the electron-hole interaction. We note that the study [29] does not provide the numerical data for the calculated exciton energies. Also, the work [28] employs the fitting method to determine the material parameters. However, the different ways of choosing the bandgap energy cause a slight discrepancy compared with our results for the parameters. The material parameters provided in Tables 2 and 2 are also consistent with our previous study [32] for monolayers WSe2 and WS2.
Although Figs. 3 (a), (b), and (c) validate our retrieved material parameters well for the case of monolayer WSe2, we also plot the magnetoexciton energies calculated with material parameters in Table 2 retrieved from the experimental data of Ref. [28] for monolayers WS2, MoSe2, and MoS2 and compare them with the experimental energies (see Figs. 4 (a), (b), and (c)).
@@
Figure 4: Magnetoexciton energies (black lines) of monolayer WS2 (a), MoSe2 (b), and MoS2 (c), calculated with the retrieved material parameters and compared with the experimental data (blue triangles) in Ref. [28].
Figure 5: Exciton binding energies calculated by numerical method with the retrieved material parameters for monolayer TMDCs in Tables 2 and 2 and compared with the experimental in Refs. [27, 29, 37, 36, 38, 28].
@@
Besides, we confirm the validity of the retrieved material parameters for the free-field case when $B=0$. Putting the complete parameters ($\mu$, $\kappa$, and $r_{0}$) into the Schrödinger equation (III) and solving it using the method given in Ref. [32], we obtain exact numerical exciton energies and compare them with the experimental ones used to retrieve the material parameters. We show the consistency between the theory and experiments in Fig. 5 for monolayer TMDCs used in Tables 2 and 2 [27, 29, 37, 36, 38, 28].

### v2

@@ III Exciton reduced mass from magnetoexciton energies
where $\alpha=r_{0}/(\kappa a_{0}^{*}){=2.088\,\xi}$; $r=\sqrt{x^{2}+y^{2}}$; energy $\varepsilon$ and coordinates $x,y$ are given in the effective Rydberg unit $\text{Ry}^{*}=\mu m_{e}e^{4}/32\pi^{2}\kappa^{2}\epsilon_{0}^{2}\hbar^{2}$ and effective Bohr radius $a_{0}^{*}=4\pi\kappa\epsilon_{0}\hbar^{2}/\mu m_{e}e^{2}$, scaled by the exciton reduced mass $\mu$ and the surrounding dielectric constant $\kappa$; $B$ is dimensionless magnetic intensity measured in the unit of $B_{0}^{*}=(2\mu m_{e}/\hbar e)\text{Ry}^{*}$; $\hbar$ and $e$ are the reduced Planck constant and electron electric charge, respectively.
Equation (III) has been derived in Refs. [33, 49], where the Laplace transform for the Rytova-Keldysh potential was used for describing the electron-hole interaction. In Ref. [49], the Schrödinger equation was first established for the whole system, an electron and a hole interacting with each other through the Rytova-Keldysh potential. After that, the center of mass motion was separated by an exact procedure; thus, the Schrödinger equation for the relative motion of an electron and a hole (an exciton with reduced mass $\mu$) was obtained. In the present paper, this equation is rewritten in a dimensionless form, where the material parameters $\mu$ and $\kappa$ consist only of scaled atomic units. The other parameter, $r_{0}$, is defined via $\alpha$, a dimensionless constant in the Laplace form of the Rytova-Keldysh potential.
For an analytical method, we need to solve the Schrödinger equation (III) to obtain exciton energies as functions of the material parameters and the magnetic intensity, given as $\varepsilon_{ns}(B,\mu,\kappa,r_{0})$. Using the relations described by Eqs. (7)-(11), we get the experimental exciton energy as a one-variable function $E_{ns}(\mu,B)$ for each value of the magnetic intensity $B$. Furthermore, we derive an inverse function $\mu(E_{ns},B)$, which allows us to determine the exciton reduced mass $\mu$ from the measured exciton energies $E_{ns}(B)$.
This kind of the Schrödinger equation (III) can be resolved analytically by the regulated perturbation method, as developed in Refs. [52, 53] for a two-dimensional exciton with the Coulomb potential for the electron-hole interaction in a magnetic field. However, in our case with the electron-hole interaction described by the Rytova-Keldysh potential, the problem is more complicated. We will find the magnetoexciton energy in the form
$\varepsilon_{ns}(B)=\varepsilon_{ns}(0)+\Delta E_{ns}(B,\mu),$ (13)
where the field-free exciton energy $\varepsilon_{ns}(0)$ is defined by Eq. (1). The magnetic part of energy $\Delta E_{ns}(B,\mu)$ is obtained from the equation as a perturbative correction in the weak regime of the magnetic field ($B\ll 1$), given that the limit of magnetic intensity is less than 120 Tesla in most current experiments.
Considering that perturbation theory, applied to the Schrödinger equation (III), provides the energy correction in the weak-field regime as a polynomial of magnetic intensity $B$, with two limits, $\sim B^{2}$ at the weak-field regime and $\sim B$ at the strong-field regime [52, 53], we suggest the correction in the Padé approximant as
$\Delta E_{ns}=\frac{\alpha_{n}B^{2}}{1+\beta_{n}B}\,\text{Ry}^{*}\,$, (14)
where coefficients $\alpha_{n}$ and $\beta_{n}$ can be calculated by the second-order perturbation theory. We suggest an alternative approach by fitting the formula (14) with numerical exact solutions of the Schrödinger equation (III). Since the energy correction $\Delta E$ weakly depends on the material parameter $\xi$, the achieved coefficients $\alpha_{n}$ and $\beta_{n}$ are actually slightly material-dependent. We provide them in Table 1 for the case of $\xi=0.4135$ (corresponding to the experimental data of Stier et al. [28]) and for $ns$ states, $n=1,2,3,4,5$.
Table 1: Coefficients $\alpha_{n}$ and $\beta_{n}$ in the formula (14) for nsns states ($n=1,2,3,4,5$) for the case of $\xi=0.4135$.
$n$ 1 2 3 4 5
$\alpha_{n}$ 0.5302 8.7698 45.2041 145.2032 217.8212
$\beta_{n}$ 0.4244 3.0909 10.1873 23.5106 32.9012
In Eq. (14), the material parameters $\mu$ and $\kappa$ are hidden in the effective atomic units for energies and magnetic intensity: $\text{Ry}^{*}$ and $B_{0}^{*}$. To have this equation in an explicit form of the exciton reduced mass $\mu$, we utilize the standard atomic units Ry and $B_{0}$ instead of the effective units for the exciton energy and magnetic intensity, using the transformations: $\text{Ry}^{*}=\mu m_{e}e^{4}/(32\pi^{2}\epsilon_{0}^{2}\kappa^{2}\hbar^{2})=(\mu/\kappa^{2})\text{Ry}$ and $B_{0}^{*}=(2\mu/e\hbar)\text{Ry}^{*}=(\mu^{2}/\kappa^{2})B_{0}$. As a consequence, we have
$\Delta E_{ns}(B,\mu)=\frac{1}{\mu}\frac{\alpha_{n}\,B^{2}}{\eta\mu+\beta_{n}B}\text{Ry}\,$, (15)
where the units of energy and magnetic intensity are now given by $\text{Ry}=m_{e}e^{4}/(32\pi^{2}\epsilon_{0}^{2}\hbar^{2})=13.60569$ eV and $B_{0}=(2m_{e}/e\hbar)\text{Ry}=235,051.7$ Tesla, which are material-independent. In Eq. (15), the value $\eta=\mu/\kappa^{2}$ is determined from the experimental field-free exciton energies by the formula (8).
Formula (15) is highly accurate for a wide class of TMDC monolayers, including WSe2, WS2, MoSe2, and MoS2 on hBN slabs. Compared with the exact numerical solutions, this formula gives an accuracy of less than 0.5 meV, so we can use it to retrieve the exciton reduced mass $\mu$ from experimental magnetoexciton energies. Indeed, given the formula for the exciton energy $E_{ns}(B)=\varepsilon_{ns}(0)+\Delta E_{ns}(B,\mu)+E_{g}$, the quantity $\delta_{B}$, defined by the following equation
$\delta_{B}$ $=$ $\frac{\Delta E_{ns}(B_{1},\mu)-\Delta E_{ns}(B_{2},\mu)}{\alpha_{n}(B_{1}-B_{2})}$ (16)
$=$ $\frac{E_{ns}(B_{1})-E_{ns}(B_{2})}{\alpha_{n}(B_{1}-B_{2})}=\delta_{B}^{exp},$
can be obtained from any two points of experimental data $E(B_{1})$ and $E(B_{2})$ at magnetic field values $B_{1}$ and $B_{2}$. Using the formula (15), we lead Eq. (16) to an equation for the exciton reduced mass $\mu$ as
$\mu^{3}+a\mu^{2}+b\mu+c=0$ (17)
with
$a$ $=$ $\frac{\beta_{n}}{\eta}(B_{1}+B_{2}),$
$b$ $=$ $\frac{{\beta_{n}}^{2}}{\eta^{2}}B_{1}B_{2}-\frac{1}{\eta\,\delta_{B}^{exp}}(B_{1}+B_{2}),$
$c$ $=$ $-\frac{\beta_{n}}{\eta^{2}\delta_{B}^{exp}}B_{1}B_{2}.$ (18)
Equation (III) is easy to solve for every two picked data points, and the final result is an average of all solutions.
We can explain the physical meaning of the coefficients $\alpha_{n}$, which are associated with the diamagnetic constants, defined by the equation:
$\sigma_{ns}$ $=$ $\frac{1}{2}\mathop{\lim}\limits_{B\to 0}\frac{\partial^{2}\Delta E_{ns}}{{\partial B}^{2}}=\frac{\alpha_{n}\kappa^{2}}{\mu^{3}}\times\frac{{\mu_{B}}^{2}}{\text{Ry}},$ (19)
where $\mu_{B}=e\hbar/2m_{e}=5.788381\times 10^{-2}$ meV/T is the Bohr magneton; and ${\mu_{B}}^{2}/{\text{Ry}}=0.24626\times 10^{-3}\,\mu{\text{eV}}/{\text{T}}$. Equation (19) can be used as an analytical formula for diamagnetic constants, which agrees excellently with the numerical calculation obtained using the traditional formulation $\sigma_{ns}=\frac{e^{2}}{8\mu m_{e}}{r_{ns}}^{2}$, where the exciton radii are calculated by equation $r_{ns}=\sqrt{\langle\psi_{ns}|r^{2}|\psi_{ns}\rangle}$, with $\psi_{ns}$ representing field-free wave functions. Using Eq. (19), we also arrive to a formula for exciton radii as
$r_{ns}$ $=$ $\frac{2\kappa}{\mu}\sqrt{\alpha_{n}}\,a_{0},$ (20)
where $a_{0}=0.0529$ nm is the Bohr radius.
@@
From the above-formulated analytical expressions, we derive a practical scheme, as shown in Fig. 1, for retrieving material parameters from experimental magnetoexciton energies.
@@
Applying this retrieval scheme to the experimental data of Refs. [28, 30, 39, 38, 40], we retrieve the material parameters for monolayer WSe2. We provide them in Table 2 and compare them with those presented in the references. More data retrieved in the same manner from Ref. [29] for other TMDC monolayers, such as WS2, MoS2, and MoSe2, are also provided in Table 3 and compared with others. Knowing the material parameters, we also calculate other essential properties, such as the diamagnetic constants and exciton radii, using equations (19) and (20). Table 4 presents the results obtained in our study, which are compared with experimental and theoretical data from references.
@@
The first remark from Table 2 is that although various experiments were performed for the same material, monolayer WSe2 encapsulated by hBN slabs, the surrounding dielectric constants $\kappa$ from these experiments differ slightly due to variations in the experimental setups. However, the exciton reduced mass $\mu$ is weakly dependent on the surrounding dielectric constant $\kappa$, as shown in Table 2. For various experiments, the value is almost $0.19\,m_{e}$, differing from the value, $0.20\,m_{e}$, used in these experiments for theoretical explanations. With the exciton reduced mass known, our method is also applicable to photoluminescence experiments without a magnetic field, as shown in Table 2 for Ref. [40]. We also note that the exciton reduced mass value in monolayer WSe2, as calculated in the present study from the experimental data, agrees better with the theoretical calculations by density functional theory (DFT) [21, 22].
The second remark is about the surrounding dielectric constant, which indeed is independent of the materials. For various experiments [28, 30, 38, 40, 29] with different TMDC monolayers (WSe2, WS2, MoSe2, and MoS2), the results are nearly identical (except in Chen et al. [39]), with values approximately equal to the bulk h-BN dielectric constant of 4.5, as referenced in [5].
@@
We also note the slight discrepancy in Tables 2 and 3 between the bandgap energies calculated by our method and those in the references. By applying the bandgap energies we calculated, the experimental binding energies for the $1s$ state (in parentheses) coincide with the prediction of our theory. The consistency of our predicted energy transitions $\Delta E_{1s-2s}$ and $\Delta E_{2s-3s}$ to the experimental data is much better than the theoretical modeling by the Keldysh potential in Refs. [28, 30, 39] or by the Kratzer potential in Ref. [38] for the electron-hole interaction. Our results are also consistent with our previous study [33] for monolayers WSe2 and WS2.
Table 2: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WSe2 retrieved from magnetoexciton energies of different experiments [28, 30, 39, 38, 40] using analytical expressions (9), (10), (11), and (17) and compared with those given in these references.
WSe2 $\xi$ $r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$
nm $(m_{e})$ meV meV eV meV
Stier et al. – exp. (169) 130 22
– theory [28] 4.50 0.20 4.50 1.890 167 124 21.3
Present work’s theory 0.4135 4.35 0.19 4.25 1.892 168.8 130.0 22.0
Liu et al. – exp. (167) 131 21
– theory [30] 5.00 0.20 3.97 1.884 172 128.3 23.3
Present work’s theory 0.3268 3.89 0.185 4.47 1.879 167.4 131.0 21.0
Chen et al. – exp. (180) 131 26
– theory [39] 4.51 0.22 4.50 1.900 170 128 23
Present work’s theory 0.8483 5.52 0.19 3.35 1.907 180.0 131.0 26.0
Molas et al. – exp. (167) 129 21.5
– theory [38] 4.50 0.20 4.50 1.873 167 128.0 21.2
Present work 0.3876 4.26 0.19 4.35 1.873 166.7 129.0 21.5
Takahashi et al. – exp. (167) 129.1 21.6
– theory [40] 3.45 0.20 5.01 1.862 164 131 19
Present work 0.3940 4.29 0.19* 4.33 1.865 167.0 129.1 21.6
Table 3: Fundamental optoelectronic material parameters (bandgap energy $E_{g}$, screening length $r_{0}=2\pi\chi_{\text{2D}}$ with 2D static polarizability $\chi_{\text{2D}}$, exciton reduced mass $\mu$, surrounding dielectric constant $\kappa$) for monolayer WS2, MoSe2, and MoS2 retrieved from magnetoexciton energies of the experiment [29] using analytical expressions (9), (10), (11), and (17) and compared with those given in this reference.
$\xi$ $r_{0}$ $\mu$ $\kappa$ $E_{g}$ $E_{b}$ $\Delta E_{1s-2s}$ $\Delta E_{2s-3s}$
nm $(m_{e})$ eV meV meV meV
WS2, Goryca et al. – exp. (181) 141 23
– theory [29] 3.40 0.175 4.35 2.238 180
Present work’s theory 0.3531 3.74 0.175 4.10 2.239 181.0 141.0 23.0
MoSe2, Goryca et al. – exp. (232) 171 33
– theory [29] 3.90 0.35 4.40 1.874 231
Present work’s theory 0.7045 4.02 0.345 4.22 1.875 231.7 171.2 32.8
MoS2, Goryca et al. – exp. (221) 170 29
– theory [29] 3.40 0.27 4.40 2.161 222
Present work’s theory 0.4276 3.37 0.275 4.43 2.160 221.3 170.0 29.0
We emphasize that the diamagnetic constants and corresponding exciton radii can be calculated analytically by Eqs. (19) and (20), which are very close to the ones calculated numerically, as shown in Table 4. For the 1s state, our theoretical results as well as the theoretical modelling of others agree well with the experimental data, except for the experiment of Ref. [39], where, however, our theoretical calculations are better than those in this reference. For the excited states, the experimental data are subject to considerable uncertainty due to the lack of data near the zero-point magnetic field $B=0$. In these cases, our calculations are relatively consistent with the experimental data and more so than those in the references.
Finally, we validate the retrieved material parameters in our study by plotting the magnetoexciton energies calculated using these parameters and comparing them with the experimental data. The results are provided in Figs. 3 and 4. Figure 3 presents our calculations compared with experimental magnetoexciton energies in monolayer WSe2 [28, 30, 39]. Figure 4 demonstrates the comparison for magnetoexciton energies in monolayers WS2, MoSe2, and MoS2 with the experiments of Ref. [29]. In these figures, theoretical data are calculated by two different methods: (i) by numerically solving the Schrödinger equation (III) and (ii) by analytical formulas (1) and (15). The figures obtained from the two methods are almost identical, demonstrating the high accuracy of our analytical solutions. The only discrepancy between the two methods is in the exciton energies at zero field, where the analytical formula (1) yields slight shifts (less than 0.5 meV) in energy compared to the exact numerical values.
[…]
$\mu$ $\kappa$ $\sigma_{1s}$ $r_{1s}$ $\sigma_{2s}$ $r_{2s}$ $\sigma_{3s}$ $r_{3s}$ $\sigma_{4s}$ $r_{4s}$
[…]
– theory (Stier [28]) 0.20 4.50 0.30 1.67 5.12 6.96 26.86 15.8
Our theory – anal. 0.19 4.25 0.344 1.723 5.687 7.008 29.315 15.911 94.164 28.517
– numer. method 0.19 4.25 0.329 1.686 5.633 6.977 29.393 15.938 94.879 28.635
[…]
– theory (Liu [30]) 0.20 3.97 0.31 1.68 4.86 6.66 24.2 14.86 76.3 26.37
Our theory – anal. 0.185 4.47 0.341 1.694 6.124 7.176 32.911 16.636 110.476 30.480
– numer. method 0.185 4.47 0.336 1.682 6.118 7.175 32.905 16.640 107.920 30.135
WSe2 – exp. 0.5 2.2 5.8 7.6 17.6 13.3
– theory (Chen [39]) 0.22 4.50 0.25 1.60 4.18 6.50 21.6 14.7
Our theory – anal. 0.19 3.35 0.348 1.733 4.916 6.516 23.128 14.133 70.390 24.656
– numerical 0.19 3.35 0.342 1.718 4.895 6.504 23.096 14.128 70.389 24.664
WS2 – exp.
– theory (Goryca [29]) 0.175 4.35 0.40 1.8
Our theory – anal. 0.175 4.10 0.369 1.712 6.257 7.055 33.407 16.301 109.296 29.486
– numerical 0.175 4.10 0.352 1.673 6.272 7.066 33.407 16.307 109.011 29.457
MoSe2 – exp.
– theory (Goryca [29]) 0.35 4.40 0.07 1.1
Our theory – anal. 0.345 4.22 0.080 1.117 1.180 4.302 5.714 9.466 17.731 16.675
– numerical 0.345 4.22 0.079 1.114 1.183 4.309 5.729 9.482 17.732 16.681
MoS2 – exp.
– theory (Goryca [29]) 0.27 4.40 0.12 1.2
Our theory – anal. 0.275 4.43 0.119 1.222 2.050 5.062 10.678 11.553 34.427 20.745
– numerical 0.275 4.43 0.120 1.227 2.051 5.065 10.680 11.558 34.429 20.752
@@
Figure 3: Magnetoexciton energies in monolayer WSe2 encapsulated by hBN slabs, calculated by the highly accurate numerical method (black lines) and by the analytical method (formulas (1) and (15)) (red lines) compared with the experimental data (blue triangles) of Refs. [28] (a), [30] (b), and [39] (c). Material parameters used in the calculations are retrieved from the experimental data using the formulas (9) (for the band gap energy $E_{g}$), (10) (for the screening length $r_{0}$), (11) (for the dielectric constant $\kappa$), and (17) (for the exciton reduced mass $\mu$).
Figure 4: Magnetoexciton energies of monolayer WS2 (a), MoSe2 (b), and MoS2 (c), calculated numerically (black lines) and analytically (red lines) with the retrieved material parameters and compared with the experimental data (blue triangles) in Ref. [29]. Material parameters used in the calculations are retrieved from the experimental data using the formulas (9) (for the band gap energy $E_{g}$), (10) (for the screening length $r_{0}$), (11) (for the dielectric constant $\kappa$), and (17) (for the exciton reduced mass $\mu$).
There is a relatively significant shift of approximately 4.0 meV between numerical and analytical calculations of the bound state in the Chen et al. experiment (Fig. 3 (c)). However, the extrapolation from the magnetoexciton energies yields an estimated value of 1,724 meV, compared to the experimental value of 1,727 meV for the $1s$-state energy, which leads to a discrepancy in calculating the material parameter. Indeed, using the extrapolated value of the bound state energy, we obtain $\xi=0.7481$, $\kappa=3.47$, $r_{0}=5.23$ nm, and $\mu=0.19$. The bandgap energy is now calculated as $E_{g}=1906.3$ meV, which leads to a bound energy of 182.32 meV. By incorporating the new parameters, our analytical and numerical calculations for the ground-state energies are much more consistent.
@@
Besides, we also confirm the validity of the retrieved material parameters for the free-field case when $B=0$. Putting the complete parameters ($\mu$, $\kappa$, and $r_{0}$) into the Schrödinger equation (III) and solving it with $B=0$ using the method given in Ref. [33], we obtain exact numerical exciton energies and compare them with the experimental ones used to retrieve the material parameters. We show the consistency between the theory and experiments in Fig. 5 for monolayer TMDCs used in Tables 2 and 3 [28, 30, 39, 38, 40, 29].
We note the boundary of our theory, which is related to the working range $0.85>\xi>0.29$ of the formula (1) for the $s$-state exciton energies. For all experiments involving monolayer TMDCs in an hBN surrounding environment, including those considered in the present study, the value $\xi$ falls in this range. However, for the experiment of Chernikov et al. [34], where the monolayer WS2 lies on a thin SiO2 layer, the screening parameter $\xi\sim 5.0$ falls outside the range. As shown in Ref. [47], the working range ensures good precision with a relative error of less than $5\%$. However, for values far outside the range, the calculation error may be significant. We will further develop the analytical formula for exciton energies to ensure its working range encompasses both free-standing and other surrounding media, including those of current interest [54].
Figure 5: Exciton binding energies calculated by numerical method with the retrieved material parameters for monolayer TMDCs in Tables 2 and 3 and compared with the experimental in Refs. [28, 30, 39, 38, 40, 29].